You come into lab and notice your lab partner did a 1:10 dilution series on a stock that you ordered. You see that the final dilution has a concentration of 0.0004 Molar and that your partner had diluted the stock seven times. What is the concentration of the original stock solution?

is it 40,0000?
Answer #1

First method:

Stock solution is diluted seven times. First time it is diluted, the dilution is 10-1 or 1: 10 dilution. Hence, final dilution is 10-7 dilution.

Final dilution= 10-7

Final concentration= 0.0004M

Initial concentration= ?= A

Initial dilution =100=1

C1D1= C2D2

A X 1= 0.0004M X 107

A= 4000M

Another method

7th dilution= 10-7 dilution = 0.0004 M

6th dilution=10-6 dilution = 0.004 M

5th dilution=10-5 dilution = 0.04 M

4th dilution=10-4 dilution = 0.4 M

3rd dilution=10-3 dilution = 4 M

2nd dilution=10-2 dilution = 40 M

1st dilution=10-1 dilution = 400 M

100 dilution = Stock solution= 4000 M

Right answer= 4000 M